OpenStudy (anonymous):

solve the equation 2a^2-18a=0

OpenStudy (raden):

first, divided by 2 from both sides : (2a^2-18a)/2 =0/2 a^2 - 9a = 0 take the common, is a then factor out! a(a - 9) = 0

OpenStudy (raden):

solve for a

OpenStudy (anonymous):

Simplifying 2a2 + -18a = 0 Reorder the terms: -18a + 2a2 = 0 Solving -18a + 2a2 = 0 Solving for variable 'a'. Factor out the Greatest Common Factor (GCF), '2a'. 2a(-9 + a) = 0 Ignore the factor 2. Subproblem 1 Set the factor 'a' equal to zero and attempt to solve: Simplifying a = 0 Solving a = 0 Move all terms containing a to the left, all other terms to the right. Simplifying a = 0 Subproblem 2 Set the factor '(-9 + a)' equal to zero and attempt to solve: Simplifying -9 + a = 0 Solving -9 + a = 0 Move all terms containing a to the left, all other terms to the right. Add '9' to each side of the equation. -9 + 9 + a = 0 + 9 Combine like terms: -9 + 9 = 0 0 + a = 0 + 9 a = 0 + 9 Combine like terms: 0 + 9 = 9 a = 9 Simplifying a = 9 Solution a = {0, 9}

OpenStudy (imtant):

2a(a-9)=0 then you will get a=0 and a=9..

OpenStudy (imtant):

2a=0,so a=0/2 will get 0.. a-9=0,bring -9 to the right become +9,so a=9..